# Bounding the degree of generic sharp transitivity

### v2

@@ Abstract.
Key words and phrases:
multiple transitivity, finite Morley rank
2020 Mathematics Subject Classification:
Primary 20B22; Secondary 20F11
@@ 1.2. The theorem
It is known—and nontrivial—that every connected group of finite Morley rank has one of the four types listed above [BBC07] and that the infinite simple ones cannot have mixed type [ABC08].
@@ 1.3. The strategy
Remark.
We wish to highlight that the $L$-hypothesis in the Theorem is only used when we are applying Proposition 1; or more to the point, removing the hypothesis from Proposition 1 would remove it from the Theorem.
Despite some effort trying to free Proposition 1 from the $L$-hypothesis, we were not able to realize it. What is needed is an extension of Corollary 2.6 from actions of $\operatorname{Alt}(n)$ on simple algebraic groups to actions on simple odd-type groups of finite Morley rank. One angle we entertained was to adapt the analysis of actions on degenerate type groups from [AW24]; there, the approach is inductive, driven by the fact that a degenerate-type group acted upon by a Klein $4$-group $K$ is generated by the centralizers of the involutions from $K$ [BBC07, Theorem 5]. But our attempts to generalize this fact to actions on odd-type groups failed to find traction.
There is also the intermediate question of trying to extend Corollary 2.6 to simple $L^{*}$-groups $G$ of odd type. The $L^{*}$-hypotheses assumes all proper definable sections of $G$ are $L$-groups, but makes no claim on $G$ itself. The case of simple $L^{*}$-groups with sufficiently high Prüfer rank has recently been reviewed and clarified by Cherlin: assuming such a $G$ has Prüfer $2$-rank at least $3$ and that there are certain internal restrictions on how $2$-tori act (the NTA2 hypothesis), then either $G$ is algebraic or contains a strongly embedded subgroup [Che24]. As such, a test case for extending Corollary 2.6 to the $L^{*}$-setting might be when $G$ has high Prüfer $2$-rank and a strongly embedded subgroup. Or, instead focusing on low Prüfer $2$-rank, one could try to analyze actions on groups of type $\textbf{CiBo}_{1}$, $\textbf{CiBo}_{2}$, or $\textbf{CiBo}_{3}$ (as defined in Section 2.3) with an initial focus on the minimal simple case.

### v3

@@ 1.2. The theorem
It is known—and nontrivial—that every connected group of finite Morley rank has one of the four types listed above [BP90, BBC07] and that the infinite simple ones cannot have mixed type [ABC08].
@@ 1.3. The strategy
1.4. Regarding the $L$-hypothesis
We would like to highlight that the $L$-hypothesis in the Theorem is only used when we are applying Proposition 1; or more to the point, removing the hypothesis from Proposition 1 would remove it from the Theorem.
Despite some effort trying to free Proposition 1 from the $L$-hypothesis, we were not able to realize it. What is needed is an extension of Corollary 2.6 from actions of $\operatorname{Alt}(n)$ on simple algebraic groups to actions on simple odd-type groups of finite Morley rank. One angle we entertained was to adapt the analysis of actions on degenerate type groups from [AW24]; there, the approach is inductive, driven by the fact that a degenerate-type group acted upon by a Klein $4$-group $K$ is generated by the centralizers of the involutions in $K$ [BBC07, Theorem 5]. Our attempt to generalize this to actions on odd-type groups failed to find traction, and this may well be a difficult problem.
Although the value is unclear, one could try replacing the $L$-hypothesis of Proposition 1 with the assumption that the “four-group generation” result holds for the simple definable sections of $H$, leading to an analysis of $V$-groups (as $K$ is already taken). We thought very briefly on this. A $V$-assumption moves the analysis of [AW24] along quite smoothly until it requires a lower bound on the corank of a proper definable subgroup (see [AW24, Corollary 2.5]), which uses that the degree of generic transitivity of any action of a connected degenerate type group is at most $1$. For odd-type groups there is no such universal bound, but something can still be said. For example, since the degree of generic transitivity of a connected solvable group is at most $2$, one obtains that the degree of generic transitivity of a connected minimal simple group is at most $3$. This may be enough to prove a statement along the following lines: if $\operatorname{Alt}(n)$, with $n$ sufficiently large, acts on a connected minimal simple $V$-group $H$ of finite Morley rank, then $n\leq\operatorname{rk}H+2$.
There is also the question of trying to extend Corollary 2.6 to simple $L^{*}$-groups $G$ of odd type. The $L^{*}$-hypotheses assumes all proper definable sections of $G$ are $L$-groups but makes no claim on $G$ itself. The case of simple $L^{*}$-groups with sufficiently high Prüfer rank has recently been reviewed and clarified by Cherlin: assuming such a $G$ has Prüfer $2$-rank at least $3$ and that there are certain internal restrictions on how $2$-tori act (the NTA2 hypothesis), then either $G$ is algebraic or contains a strongly embedded subgroup [Che24]. As such, a test case for extending Corollary 2.6 to the $L^{*}$-setting might be when $G$ has high Prüfer $2$-rank and a strongly embedded subgroup. Or, instead focusing on low Prüfer $2$-rank, one could try to analyze actions on groups of type $\textbf{CiBo}_{1}$, $\textbf{CiBo}_{2}$, or $\textbf{CiBo}_{3}$ (as defined in Section 2.3) with an initial focus on the minimal simple case.